3(2s-3)=-(7-s) can you show me the check too
Answer by edjones(8007)

3(2s-3)=-(7-s)
6s-9=-7+s
subtract s and add 9 to each side. 5s=2
divide 5 into each side. s=2/5=.4
Check:
3(2*.4-3)=-(7-.4)
3(.8-3)=-7+.4
2.4-9=-6.6
6.6=6.6
Ed